# Efficient Learning for Entropy-regularized Markov Decision Processes via Multilevel Monte Carlo

### v1

@@ 1 Introduction
In this work, we focus on Monte Carlo (MC) sampling algorithms, which are a popular class of methods for estimating optimal value functions by computing empirical averages over sampled trajectories. Various MC sampling algorithms have been proposed for MDPs with finite action spaces and arbitrary state spaces, achieving quasi-polynomial or polynomial sample complexity in terms of the desired accuracy (see e.g., [24, 15, 3]). However, these sample complexity guarantees depend explicitly on the cardinality of the action space and grow unbounded for large (particularly continuous) action spaces; see Section 1.2 for more details.
This work addresses this gap in the context of entropy-regularized MDPs, where the objective is augmented with an entropy term. Unlike prior works such as [15], which only consider finite action spaces, we allow both the state and action spaces to be general Polish spaces. Our key observation is that the Bellman operator of an entropy-regularized MDP involves integration over the action space. Leveraging this insight, we propose several MC algorithms that achieve provable quasi-polynomial or even polynomial sample complexity guarantees that are independent of the dimensions or cardinalities of the state and action spaces.
@@ 1.1 Outline of Main Results
Consider an infinite horizon MDP $(\mathcal{S},\mathcal{A},P,c,\gamma)$, where the state space $\mathcal{S}$ and action space $\mathcal{A}$ are Polish (i.e., complete separable metric) spaces with possibly infinite cardinality, $P\in\mathcal{P}(\mathcal{S}|\mathcal{S}\times\mathcal{A})$ is the transition probability kernel, $c$ is a bounded cost function, and $\gamma\in[0,1)$ is the discount factor. Let $\mu\in\mathcal{\mathcal{P}}(\mathcal{A})$ denote a reference probability measure and $\tau>0$ denote a regularisation parameter. For each stochastic policy $\pi\in\mathcal{P}(\mathcal{A}|\mathcal{S})$ and $s\in\mathcal{S}$, define the regularised value function by
@@
By the dynamic programming principle [26, Appendix B], both the optimal function $V^{\star}$ and the optimal policy that minimizes (1.1) are given by
$V^{\star}(s)=-\tau\log\int_{\mathcal{A}}\exp\left(-\frac{1}{\tau}Q^{\star}(s,a)\right)\mu({\mathrm{d}}a),\quad\pi^{\star}({\mathrm{d}}a|s)=\exp\left(-\frac{Q^{\star}(s,a)-V^{\star}(s)}{\tau}\right)\mu({\mathrm{d}}a),$ (1.2)
@@
$TQ(s^{\prime})=-\tau\log\int_{\mathcal{A}}\exp\left(-\frac{1}{\tau}Q(s^{\prime},a^{\prime})\right)\mu({\mathrm{d}}a^{\prime}),$ (1.4)
@@
In this paper, motivated by the identity (1.2), we propose and analyze several MC estimators for $Q^{\star}$, utilizing an oracle that generates state transition samples from arbitrary state-action pairs and evaluates the corresponding instantaneous cost $c$, along with a sampler for the reference measure $\mu$.
@@
${Q}_{n,M,K}(s,a)$ $\coloneqq c(s,a)+\frac{1}{M}\sum_{i=1}^{M}\hat{T}{Q}_{n-1,M,K}(S_{s,a}^{(n-1,i)}),$ (1.6)
$\hat{T}{Q}_{n-1,M,K}(s)$ $\coloneqq-\tau\log\frac{1}{K}\sum_{k=1}^{K}\exp\left(-\frac{Q(s,A^{(n-1,k)})}{\tau}\right),$
@@
The first main contribution of this work is to analyze the sample complexity of the estimator (1.6). In particular,
  • We explicitly quantify the $L^{2}$ error of the estimator ${Q}_{n,M,\mathbf{T}_{K}}$ in terms of $M$, $K$ and $n$ (Theorem 2.1). Leveraging this error bound, we prove the estimator (1.6) achieves accuracy $\varepsilon$ with a quasi-polynomial complexity of the order $\varepsilon^{-\kappa\log\varepsilon}$ as $\varepsilon\to 0$, for some $\kappa>0$. Note that this error bound is independent of the cardinality of the action space, in contrast to the quasi-polynomial complexity bound in [24].
The error estimate in Theorem 2.1 also indicates that the estimator (1.6) cannot achieve polynomial sample complexity. This is due to the $\mathcal{O}(M^{-1/2})$ approximation error for the expectation over $\mathcal{S}$ at each iteration, resulting in an overall sample complexity of at least $\mathcal{O}(M^{n})$ (Remark 2.3). This motivates us to adopt the multilevel Monte Carlo (MLMC) technique, originally proposed in [13], to achieve variance reduction.
@@
$\quad+\gamma\sum_{l=1}^{n-1}\sum_{i=1}^{M^{n-l}}\frac{1}{M^{n-l}}\left[\hat{T}Q_{l,M}(S^{(l,i)}_{s,a})-\hat{T}Q_{l-1,M}(S^{(l,i)}_{s,a})\right],$
@@
The MLMC estimator (1.7) differs from the estimator in [3], which was developed specifically for unregularized MDPs with finite action spaces. The key distinction is that (1.7) allows for a more general class of stochastic operators $\hat{T}$ to approximate the soft-Bellman operator $T$, a crucial feature for constructing an MLMC estimator that can accommodate general action spaces. In contrast, [3] fixes $\hat{T}$ as the (exact) Bellman operator for the unregularized MDP, which requires evaluating a given Q-function at all actions and taking the maximum over them. This approach does not scale well to large action spaces and is inapplicable to our setting with general action spaces.
@@
    The first choice of $\hat{T}$ is the plain MC estimator (1.6), which serves as a biased approximation of the soft-Bellman operator $T$ due to the logarithm function in $T$. We prove that the corresponding MLMC estimator (1.7) achieves a quadratic reduction in sample complexity compared to the simple iterative MC estimator (1.6), highlighting the advantage of the MLMC technique (Theorem 2.3). However, the inherent bias in $\hat{T}$ causes the overall complexity to remain quasi-polynomial (Remark 2.6).
@@
Plain MC (iterative) Theorem 2.1 $\mathcal{O}\left(\frac{1}{\sqrt{M}}+\frac{1}{\sqrt{K}}+(\gamma L)^{n}\right)$ $\varepsilon^{\frac{-4\log\varepsilon}{\log\gamma L}(1+o(1))}$
MLMC (biased) Theorem 2.3 $\mathcal{O}\left((\Lambda_{M})^{n}+\frac{1}{\sqrt{K}}\right)$ $\varepsilon^{\frac{-2\log\varepsilon}{\log\gamma L+\delta}(1+o(1))}$
@@ 1.2 Most Related Works
Entropy regularizations have emerged as powerful tools in RL, offering significant benefits across various aspects of algorithm design and performance [11]. These techniques are known to stabilize learning [43, 30] and prevent the agent from being trapped in suboptimal policies too early [10]. This approach has given rise to popular deep RL methods such as soft actor-critic (SAC) [18] and proximal policy optimization (PPO) [34], which have become staples in modern RL applications. Such regularizations also facilitate the design and study of RL algorithms in continuous time (see e.g., [21, 38]), as well as in the multi-agent/mean-field context [7, 2, 17]. The study of entropy-regularized infinite-horizon MDPs with general action and state spaces has led to important theoretical advancements, particularly in proving the convergence of policy gradient techniques [6, 26]. Our estimator shares similarities with techniques used in distributionally robust Q-learning [27, 40, 41], where suitable unbiased estimators are employed to improve state-of-the-art complexity in the tabular case [41].
@@
Several existing works utilize MLMC for the numerical approximation of equations with an underlying fixed-point structure. Multilevel Picard iterations have proven effective in overcoming the curse of dimensionality when solving semilinear PDEs [9] through the Bismuth–Elworthy–Li and Feynman–Kac formulas. Similarly, [3] introduced multilevel fixed-point iterations for learning the optimal Q-function in MDPs with a finite action space, obtaining a complexity bound that explicitly depends on the cardinality of the action space.
@@
Our problem also differs from the work of [36], which applies MLMC to estimate nested expectations with finite depth. Note that we aim to estimate the fixed point of (1.3), which cannot be expressed as a nested expectation of finite depth. More importantly, our estimator requires selecting the depth (corresponding to the level $n$ in (1.7)) of nested expectations as a function of the error $\varepsilon$, and we demonstrate that our estimator remains polynomial in $\varepsilon^{-1}$. In contrast, [36] provide a polynomial complexity result depending on the fixed depth. At the time of writing, we are not aware of any other model-free reinforcement learning techniques capable of achieving average polynomial complexity in arbitrary state and action spaces without structural assumptions on the underlying MDP.
Finally, we would like to emphasize a key difference between our MLMC estimator and Q-learning [42]. In Q-learning, the Q-values for all state-action pairs are stored (either in a look-up table for the tabular setting or using function approximation in the continuous setting), and they are updated iteratively, typically using one sample transition at a time. The convergence guarantee of Q-learning typically requires finite state and action spaces. In contrast, our estimators compute the Q-value for a specific state-action pair and require sampling multiple transitions from the oracle starting from that pair. Our convergence results hold for MDPs with general state and action spaces.
@@ 2.1 Formulation of Regularized MDPs
$\mathcal{S}$ and $\mathcal{A}$ are Polish spaces (i.e. complete separable metric spaces), $P\in\mathcal{P}(\mathcal{S}|\mathcal{S}\times\mathcal{A})$, $c\in B_{b}(\mathcal{S}\times\mathcal{A})$, $\gamma\in[0,1)$, $\mu\in\mathcal{P}(A)$ and $\tau>0$. Let $c_{\min},c_{\max}\in[0,\infty)$ be such that $c_{\min}\leq c(s,a)\leq c_{\max}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, and define $\alpha\coloneqq c_{\min}/(1-\gamma)$ and $\beta\coloneqq c_{\max}/(1-\gamma)$.
@@
In this paper, we construct MC estimators for the optimal Q-function using sampled states and actions. To this end, let $(\Omega,\mathcal{F},\mathbb{P})$ be a generic probability space that supports all (countably many) independent random variables used in the estimator, and let $\Theta=\bigcup_{n\in{\mathbb{N}}}{\mathbb{Z}}^{n}$ be the index set for these independent random variables. Note that although the practical implementation of our MLMC estimator involves only finitely many random variables, we define the estimators for all $\theta\in\Theta$ through an induction process for mathematical convenience (see Definition 2.3).
We assume access to an oracle that generates independent samples from the reference measure $\mu$ and the transition kernel $P$. To ensure the conditional independence of samples from different oracle queries, we recall the “noise outsourcing” lemma [23, Lemma 2.22]: given the kernel $P\in\mathcal{P}(\mathcal{S}|\mathcal{S}\times\mathcal{A})$, there exists a measurable function $f:\mathcal{S}\times\mathcal{A}\times[0,1]\to\mathcal{S}$ such that if $U$ is a uniform random variable on $[0,1]$, $f(s,a,U)$ has distribution $P(\cdot|s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$.
@@ 2.2 A Simple Iterative MC Estimator and Its Sample Complexity
${Q}_{n,M,\mathbf{T}_{K}}^{\theta}(s,a)$ $=c(s,a)+\frac{1}{M}\sum_{i=1}^{M}T^{(\theta,i)}{Q}_{n-1,M,\mathbf{T}_{K}}^{(\theta,i)}(S_{s,a}^{(\theta,i)}),\quad\forall n\geq 1.$
@@
and define the sample complexity $\mathfrak{C}_{n,M,K}$ of the estimator ${Q}_{n,M,\mathbf{T}_{K}}^{\theta}$ as the total number of random variables required to evaluate ${Q}_{n,M,\mathbf{T}_{K}}^{\theta}$. Notice that $E_{n,M,K}$ is independent of $\theta$ since $\left(Q^{\theta}_{n,M,\mathbf{T}_{K}}(s,a)-Q^{\star}(s,a)\right)_{\theta\in\Theta}$ are identically distributed.
@@
$E_{n,M,K}\leq\frac{\gamma\sqrt{C}}{\sqrt{M}(1-\gamma L)}+\frac{\gamma L^{\prime}}{\sqrt{K}(1-\gamma L)}+(\gamma L)^{n}\left\|Q_{0}-Q^{\star}\right\|_{\infty},$ (2.5)
with $C=\left(\beta-\alpha\right)^{2}$ and $L^{\prime}=\tau\exp\left(\tau^{-1}\beta\right)\left[\exp\left(\beta\right)-\exp\left(\alpha\right)\right]$. Moreover, the corresponding sample complexity $\mathfrak{C}_{n,M,K}$ of $Q^{\theta}_{n,M,\mathbf{T}_{K}}$ is $M^{n}K^{n}$.
@@
$n_{\varepsilon}=\left\lceil\frac{\log\varepsilon-\log(3\left\|Q_{0}-Q^{\star}\right\|_{\infty})}{\log\gamma L}\right\rceil,\quad M_{\varepsilon}=\left\lceil\frac{9\gamma^{2}C}{(1-\gamma L)^{2}\varepsilon^{2}}\right\rceil,\quad K_{\varepsilon}=\left\lceil\frac{9\gamma^{2}(L^{\prime})^{2}}{(1-\gamma L)^{2}\varepsilon^{2}}\right\rceil,$ (2.6)
it holds that $E_{n_{\varepsilon},M_{\varepsilon},K_{\varepsilon}}\leq\varepsilon$ for all $\varepsilon\in(0,1)$, and $\mathfrak{C}_{n_{\varepsilon},M_{\varepsilon},K_{\varepsilon}}=\varepsilon^{\frac{-4\log\varepsilon}{\log\gamma L}(1+o(1))}$ as $\varepsilon\to 0$, where $o(1)$ denotes a term that vanishes as $\varepsilon\to 0$.
@@
The extra assumption $\gamma L<1$ made in Theorem 2.1 holds for a sufficiently small discount factor $\gamma$, a sufficiently flat cost $c$, or a sufficiently large regularization parameter $\tau$. Indeed, for any given bounded cost $c$ and regularisation parameter $\tau$, it is satisfied if the discount factor $\gamma$ is sufficiently small. Conversely, for any $\gamma<1$, it is satisfied if either $\tau$ is sufficiently large for given $c$, or $c$ is sufficiently flat (i.e., $c_{\max}-c_{\min}$ sufficiently small) for given $\tau$.
Remark 2.3 (Error decompositions).
The error bound in (2.5) quantifies the contributions of three distinct error sources. The first term represents the variance associated with approximating the expectation over the state space $\mathcal{S}$, the second term accounts for the bias in approximating the soft-Bellman operator $T$, while the third term reflects the error introduced by the fixed-point iteration.
Theorem 2.1 indicates that the simple iterative estimator ${Q}_{n,M,\mathbf{T}}^{\theta}$ cannot achieve a polynomial sample complexity, even if the soft-Bellman operator $T$ can be evaluated exactly. This is due to the $\mathcal{O}(M^{-1/2})$ approximation error for the expectation over $\mathcal{S}$ at each iteration, which yields a sample complexity of at least $\mathcal{O}(M^{n})$. Since both $M$ and $n$ must increase to achieve a higher accuracy, the simple iterative estimator exhibits super-polynomial complexity.

### v3

@@ 1 Introduction
In this work, we focus on Monte Carlo (MC) sampling algorithms, which are a popular class of methods for estimating optimal value functions by computing empirical averages over sampled trajectories. Various MC sampling algorithms have been proposed for MDPs with finite action spaces and arbitrary state spaces, achieving quasi-polynomial or polynomial sample complexity in terms of the desired accuracy (see, e.g., [25, 15, 3]). However, these sample complexity guarantees depend explicitly on the cardinality of the action space and grow unbounded for large (particularly continuous) action spaces; see Section 1.2 for more details.
This work addresses this gap in the context of entropy-regularized MDPs, where the objective is augmented with an entropy term. Unlike prior works such as the paper of [15], which only considers finite action spaces, we allow both the state and action spaces to be general Polish spaces. Our key observation is that the Bellman operator of an entropy-regularized MDP involves integration over the action space. Leveraging this insight, we propose several MC algorithms that achieve provable quasi-polynomial or even polynomial sample complexity guarantees that are independent of the dimensions or cardinalities of the state and action spaces.
@@ 1.1 Outline of Main Results
Consider an infinite horizon MDP $(\mathcal{S},\mathcal{A},P,c,\gamma)$, where the state space $\mathcal{S}$ and action space $\mathcal{A}$ are Polish (i.e., complete separable metric) spaces with possibly infinite cardinality, $P\in\mathcal{P}(\mathcal{S}|\mathcal{S}\times\mathcal{A})$ is the transition probability kernel, $c$ is a bounded cost function, and $\gamma\in[0,1)$ is the discount factor. Let $\mu\in\mathcal{\mathcal{P}}(\mathcal{A})$ denote a reference probability measure and $\tau>0$ denote a regularization parameter. For each stochastic policy $\pi\in\mathcal{P}(\mathcal{A}|\mathcal{S})$ and $s\in\mathcal{S}$, define the regularized value function by [^1] The fact that $\mathcal{S}$ and $\mathcal{A}$ are Polish spaces allows for applying the Kolmogorov extension theorem to construct the unique probability measure associated with the kernel $P$ and a policy $\pi$, ensuring that (1.1) is well-defined; see [27, Section 2.1].
@@
By the dynamic programming principle [27, Appendix B], both the optimal function $V^{\star}$ and the optimal policy that minimizes (1.1) are given by
$V^{\star}(s)=-\tau\log\int_{\mathcal{A}}\exp\left(-\frac{Q^{\star}(s,a)}{\tau}\right)\mu({\mathrm{d}}a),\quad\pi^{\star}({\mathrm{d}}a|s)=\exp\left(-\frac{Q^{\star}(s,a)-V^{\star}(s)}{\tau}\right)\mu({\mathrm{d}}a),$ (1.2)
@@
$TQ(s^{\prime})=-\tau\log\int_{\mathcal{A}}\exp\left(-\frac{Q(s^{\prime},a^{\prime})}{\tau}\right)\mu({\mathrm{d}}a^{\prime}),$ (1.4)
@@
In this paper, motivated by the identity (1.2), we propose and analyze several MC estimators for $Q^{\star}$, using an oracle that generates state transition samples from arbitrary state-action pairs and evaluates the corresponding instantaneous cost $c$, along with a sampler for the reference measure $\mu$.
@@
${Q}_{n,M,K}(s,a)$ $\coloneqq c(s,a)+\frac{\gamma}{M}\sum_{i=1}^{M}\hat{T}{Q}_{n-1,M,K}\left(S_{s,a}^{(n-1,i)}\right),$ (1.6)
$\hat{T}{Q}_{n-1,M,K}(s)$ $\coloneqq-\tau\log\frac{1}{K}\sum_{k=1}^{K}\exp\left(-\frac{Q_{n-1,M,K}\left(s,A^{(n-1,k)}\right)}{\tau}\right),$
@@
The first main contribution of this work is to analyze the sample complexity of the estimator defined in (1.6). In particular,
  • We explicitly quantify the $L^{2}$ error of the estimator ${Q}_{n,M,\mathbf{T}_{K}}$ in terms of $M$, $K$ and $n$ (Theorem 2.1). Leveraging this error bound, we prove the estimator (1.6) achieves accuracy $\varepsilon$ with a quasi-polynomial complexity of the order $\varepsilon^{-\kappa\log\varepsilon}$ as $\varepsilon\to 0$, for some $\kappa>0$. This error bound is independent of the cardinality of the action space, in contrast to the quasi-polynomial complexity bound in [25].
The error estimate in Theorem 2.1 also indicates that the estimator (1.6) cannot achieve polynomial sample complexity. This is due to the $\mathcal{O}(M^{-1/2})$ approximation error for the expectation over $\mathcal{S}$ at each iteration, resulting in an overall sample complexity of at least $\mathcal{O}(M^{n})$ (Remark 2.3). This motivates us to adopt the multilevel Monte Carlo (MLMC) technique, originally proposed by [13], to achieve variance reduction.
@@
$\quad+\gamma\sum_{l=1}^{n-1}\sum_{i=1}^{M^{n-l}}\frac{1}{M^{n-l}}\left[\hat{T}Q_{l,M}\left(S^{(l,i)}_{s,a}\right)-\hat{T}Q_{l-1,M}\left(S^{(l,i)}_{s,a}\right)\right],$
@@
The MLMC estimator (1.7) differs from the estimator in [3], which was developed specifically for unregularized MDPs with finite action spaces. The key distinction is that (1.7) allows for a general class of stochastic operators $\hat{T}$ to approximate the soft-Bellman operator $T$, a crucial feature for constructing an MLMC estimator that can accommodate general action spaces. In contrast, [3] fix $\hat{T}$ as the (exact) Bellman operator for the unregularized MDP, which requires evaluating a given Q-function at all actions and taking the maximum over them. This approach does not scale well to large action spaces and is inapplicable to our setting with general action spaces.
@@
    The first choice of $\hat{T}$ is the plain MC estimator (1.6), which serves as a biased approximation of the soft-Bellman operator $T$ due to the logarithm function in $T$ (1.4). We prove that the corresponding MLMC estimator (1.7) achieves a cubic reduction in sample complexity compared to the simple iterative MC estimator (1.6), highlighting the advantage of the MLMC technique (Theorem 2.3). However, the inherent bias in $\hat{T}$ causes the overall complexity to remain quasi-polynomial (Remark 2.6).
@@
Plain MC (iterative) Theorem 2.1 $\mathcal{O}\left(\frac{1}{\sqrt{M}}+\frac{1}{K}+(\gamma L)^{n}\right)$ $\varepsilon^{\frac{-3\log\varepsilon}{\log\gamma L}(1+o(1))}$
MLMC (biased) Theorem 2.3 $\mathcal{O}\left((\Lambda_{M})^{n}+\frac{1}{K}\right)$ $\varepsilon^{\frac{-\log\varepsilon}{\log\gamma L+\delta}(1+o(1))}$
@@ 1.2 Most Related Works
Entropy regularizations have emerged as powerful tools in RL, offering significant benefits across various aspects of algorithm design and performance [11]. These techniques are known to stabilize learning [45, 31] and prevent the agent from being trapped in suboptimal policies too early [10]. This approach has given rise to popular deep RL methods such as soft actor-critic (SAC) [18] and proximal policy optimization (PPO) [35], which have become staples in modern RL applications. Such regularizations also facilitate the design and study of RL algorithms in continuous time (see, e.g., [22, 40]), as well as in the multi-agent/mean-field context [7, 2, 17]. The study of entropy-regularized infinite-horizon MDPs with general action and state spaces has led to important theoretical advancements, particularly in proving the convergence of policy gradient techniques [6, 27]. Our estimator shares similarities with techniques used in distributionally robust Q-learning [28, 42, 43], where suitable unbiased estimators are employed to improve state-of-the-art complexity in the tabular case [43].
@@
Iterative MLMC has been applied to approximate nonlinear equations with an underlying fixed-point structure (see, e.g., [20, 14, 9] and the references therein for applications to PDEs, as well as [39, 21] for applications to McKean-Vlasov SDEs). Recently, [3] introduced multilevel fixed-point iterations for learning the optimal Q-function of unregularized MDPs with a finite action space, obtaining a polynomial complexity bound that explicitly depends on the cardinality of the action space.
@@
Our problem also differs from the work of [37], which applies MLMC to estimate nested expectations with finite depth. Note that we aim to estimate the fixed point of (1.3), which cannot be expressed as a nested expectation of finite depth. More importantly, our estimator requires selecting the depth (corresponding to the level $n$ in (1.7)) of nested expectations as a function of the error $\varepsilon$, and we demonstrate that our estimator remains polynomial in $\varepsilon^{-1}$. In contrast, [37] provide a polynomial complexity result, depending exponentially on the fixed depth. At the time of writing, we are not aware of any other model-free reinforcement learning techniques capable of achieving average polynomial complexity in arbitrary state and action spaces without structural assumptions on the underlying MDP.
Finally, we would like to point out the difference between our MLMC estimator and Q-learning [44]. In Q-learning, the Q-values for all state-action pairs are stored (either in a look-up table for the tabular setting or using function approximation in the continuous setting), and they are updated iteratively, typically using one sample transition at a time. The convergence guarantee of Q-learning typically requires finite state and action spaces. In contrast, our estimators compute the Q-value for a specific state-action pair and require sampling multiple transitions from the oracle starting from that pair. Our convergence results hold for MDPs with general state and action spaces.
@@ 2.1 Formulation of Regularized MDPs
$\mathcal{S}$ and $\mathcal{A}$ are Polish spaces (i.e.,  complete separable metric spaces), $P\in\mathcal{P}(\mathcal{S}|\mathcal{S}\times\mathcal{A})$, $c\in B_{b}(\mathcal{S}\times\mathcal{A})$, $\gamma\in[0,1)$, $\mu\in\mathcal{P}(A)$ and $\tau>0$. Let $c_{\min},c_{\max}\in[0,\infty)$ be such that $c_{\min}\leq c(s,a)\leq c_{\max}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, and define $\alpha\coloneqq c_{\min}/(1-\gamma)$ and $\beta\coloneqq c_{\max}/(1-\gamma)$.
@@
In this paper, we construct MC estimators for the optimal Q-function using sampled states and actions. To this end, let $(\Omega,\mathcal{F},\mathbb{P})$ be a generic probability space that supports all (countably many) independent random variables used in the estimator, and let $\Theta=\bigcup_{n\in{\mathbb{N}}}{\mathbb{Z}}^{n}$ be the index set for these independent random variables. Note that although the practical implementation of our MLMC estimator involves only finitely many random variables, we define the estimators for all $\theta\in\Theta$ through an induction process for mathematical convenience (see (2.4) and Definition 2.3).
We assume access to an oracle that generates independent samples from the reference measure $\mu$ and the transition kernel $P$. To ensure the conditional independence of samples from different oracle queries, we recall the “noise outsourcing” lemma [24, Lemma 2.22]: given the kernel $P\in\mathcal{P}(\mathcal{S}|\mathcal{S}\times\mathcal{A})$, there exists a measurable function $f:\mathcal{S}\times\mathcal{A}\times[0,1]\to\mathcal{S}$ such that if $U$ is a uniform random variable on $[0,1]$, $f(s,a,U)$ has distribution $P(\cdot|s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$.
@@ 2.2 A Simple Iterative MC Estimator and Its Sample Complexity
${Q}_{n,M,\mathbf{T}_{K}}^{\theta}(s,a)$ $=c(s,a)+\frac{\gamma}{M}\sum_{i=1}^{M}T^{(\theta,i)}{Q}_{n-1,M,\mathbf{T}_{K}}^{(\theta,i)}\left(S_{s,a}^{(\theta,i)}\right),\quad\forall n\geq 1.$
@@
and define the sample complexity $\mathfrak{C}_{n,M,K}$ of the estimator ${Q}_{n,M,\mathbf{T}_{K}}^{\theta}$ as the total number of random variables required to evaluate ${Q}_{n,M,\mathbf{T}_{K}}^{\theta}$. Notice that $E_{n,M,K}$ is independent of $\theta$ since $(Q^{\theta}_{n,M,\mathbf{T}_{K}}(s,a)-Q^{\star}(s,a))_{\theta\in\Theta}$ are identically distributed.
@@
$E_{n,M,K}\leq\frac{\gamma\sqrt{C}}{\sqrt{M}(1-\gamma L)}+\frac{\gamma(L^{\prime})^{2}}{2\tau K(1-\gamma L)}+(\gamma L)^{n}\left\|Q_{0}-Q^{\star}\right\|_{\infty},$ (2.5)
with $C=\left(\beta-\alpha\right)^{2}$ and $L^{\prime}=\tau(L-1)$. Moreover, the corresponding sample complexity $\mathfrak{C}_{n,M,K}$ of $Q^{\theta}_{n,M,\mathbf{T}_{K}}$ is $M^{n}K^{n}$.
@@
$n_{\varepsilon}=\left\lceil\frac{\log\varepsilon-\log(3\left\|Q_{0}-Q^{\star}\right\|_{\infty})}{\log\gamma L}\right\rceil,\quad M_{\varepsilon}=\left\lceil\frac{9\gamma^{2}C}{(1-\gamma L)^{2}\varepsilon^{2}}\right\rceil,\quad K_{\varepsilon}=\left\lceil\frac{3\gamma(L^{\prime})^{2}}{2\tau(1-\gamma L)\varepsilon}\right\rceil,$ (2.6)
it holds that $E_{n_{\varepsilon},M_{\varepsilon},K_{\varepsilon}}\leq\varepsilon$ for all $\varepsilon\in(0,1)$, and $\mathfrak{C}_{n_{\varepsilon},M_{\varepsilon},K_{\varepsilon}}=\varepsilon^{\frac{-3\log\varepsilon}{\log\gamma L}(1+o(1))}$ as $\varepsilon\to 0$, where $o(1)$ denotes a term that vanishes as $\varepsilon\to 0$.
@@
The extra assumption $\gamma L<1$ made in Theorem 2.1 holds for a sufficiently small discount factor $\gamma$, a sufficiently flat cost $c$, or a sufficiently large regularization parameter $\tau$. Indeed, for any given bounded cost $c$ and regularization parameter $\tau$, it is satisfied if the discount factor $\gamma$ is sufficiently small. Conversely, for any $\gamma<1$, it is satisfied if either $\tau$ is sufficiently large for given $c$, or $c$ is sufficiently flat (i.e., $c_{\max}-c_{\min}$ sufficiently small) for a given $\tau$.
Remark 2.3 (Error decomposition).
The error bound in (2.5) quantifies the contributions of three distinct error sources. The first term represents the variance associated with approximating the expectation over the state space $\mathcal{S}$, the second term accounts for the bias in approximating the soft-Bellman operator $T$, and the third term reflects the error introduced by the fixed-point iteration.
Theorem 2.1 indicates that the simple iterative estimator ${Q}_{n,M,\mathbf{T}}^{\theta}$ cannot achieve a polynomial sample complexity, even if the soft-Bellman operator $T$ can be evaluated exactly. This is due to the $\mathcal{O}(M^{-1/2})$ approximation error for the expectation over $\mathcal{S}$ at each iteration combined with a sample complexity of at least $\mathcal{O}(M^{n})$. Since both $M$ and $n$ must increase to achieve a higher accuracy, the simple iterative estimator exhibits super-polynomial complexity.